OpenStudy (anonymous):

At 10am the temp was 71 degrees.At 3pm the tempature was 86 degrees. find the value of the slope and interpert in words.

OpenStudy (amistre64):

slope equals the change in (y) over the change in (x) here y=degrees and x=time(hours). (86-71)/5 hours = 15/5 = 3/1. The slope is equal to (3/1) and that means the for every hour that has passed it got 3 degrees warmer.